# Dynamic Post-Hoc Neural Ensemblers

### v1

@@ 4 Experiments and Results
4.1 Proof-of-Concept: Ensemble of Quadratic Functions
To further highlight the importance of the dynamical ensemble with diverse base models, we propose a simple regression problem with a third-degree polynomial with the ground truth function $y_{\mathrm{true}}=1.33x^{3}-0.77x^{2}-0.31x-1$, where $x\in\left[-1,1\right]$. We fit three second-degree polynomials on different subsets of the training dataset, to obtain three base functions $z_{1}(x),z_{2}(x),z_{3}(x)$. As ensemblers we consider i) a static model average $\sum_{i}\theta_{m}z_{m}(x)$, ii) a dynamic model stacker $\phi(z(x);\beta)$ and iii) model average $\sum_{m}\theta_{m}(z(x);\beta)\cdot z_{m}(x)$. In this experiment, $\phi(\cdot)$ and $\theta_{m}(\cdot)$ are two-layer MLPs with 10 hidden neurons with parameters $\beta$ and trained on the validation data. Figure 3 shows the specific data, base functions, and learned ensemblers. The rightmost plot shows that the dynamic model averaging (MSE=$0.0101$) and stacker (MSE=$0.0055$), by expressing more complex functions, better model the ground truth than the static ensemble (MSE=$0.1662$).
Figure 3: Intuition Example for Dynamical Ensembling. When considering three base models $z_{1}(x),z_{2}(x),z_{3}(x)$, the dynamical ensemblers $\phi(\cdot)$ and $\theta_{m}(\cdot)$ depending on $x$ achieve better performance than a static model average, as observable in the right-most figure.
@@ 4.2 Experimental Setup
In our experiments, we utilize four meta-datasets with pre-computed predictions, which allows us to simulate ensembles without the need to fit models. These meta-datasets cover diverse data modalities, including Computer Vision, Tabular Data, and Natural Language Processing. Additionally, we evaluate the method on datasets without pre-computed predictions to assess the performance of ensembling methods that require model fitting. Table 1 reports the main information related to these datasets. Particularly for Nasbench, we created 2 versions by subsampling 100 and 1000 models. The metadataset for Finetuning Text Classifiers (FTC) was generated by ourselves to evaluate ensembling techniques on text classification tasks by finetuning language models such as GPT2 (Radford et al., 2019), Bert (Devlin et al., 2018) and Bart (Lewis et al., 2019). We also generate a set of fitted Scikit-Learn Pipelines on classification datasets. In this case, we stored the pipeline in memory, allowing us to evaluate our method in practical scenarios where the user has fitted models instead of predictions. We detailed information about the creation of these two metadatasets in Appendix D. Information about each dataset lies in the respective referred work under the column Task Information in Table 1.
Table 1: Metadatasets Information.
Meta-Dataset Modality Task Information No. Datasets Avg. Samples for Validation Avg. Samples for Test Avg. Models per Dataset
Nasbench (100) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 100
Nasbench (1K) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 1K
QuickTune (Micro) Vision Finetuning, Classification (Arango et al., 2024) 30 160 160 255
QuickTune (Mini) Vision Finetuning, Classification (Arango et al., 2024) 30 1088 1088 203
FTC Language Finetuning, Classification, Section D.1 6 39751 29957 105
TabRepo Clas. Tabular Classification  (Salinas & Erickson, 2023) 83 1134 126 1530
TabRepo Reg. Tabular Regression  (Salinas & Erickson, 2023) 17 3054 3397 1530
Sk-Learn Pipelines. Tabular Classification, Section D.2 69 1514 1514 500
@@
We compare the Neural Ensemblers (NE) with other common and competitive ensemble approaches. 1) Single best selects the best model according to the validation metric; 2) Random chooses randomly $N=50$ models to ensemble, 3) Top-N ensembles the best $N$ models according to the validation metric; 4) Greedy creates an ensemble with $N=50$ models by iterative selecting the one that improves the metric as proposed by previous work (Caruana et al., 2004); 5) Quick builds the ensemble with 50 models by adding model subsequently only if they strictly improve the metric; 6) CMAES (Purucker & Beel, 2023b) uses an evolutive strategy with a post-processing method for ensembling, 7) Model Average (MA) computed the sum of the predictions with constant weights as in Equation 3. We also compare to methods that perform ensemble search iteratively via Bayesian Optimization such as 8) DivBO (Shen et al., 2022), and 9) Ensemble Optimization (EO) (Levesque et al., 2016). Finally, we report results by using common ML models as stackers, such as 10) SVM, 11) Random Forest, 12) Gradient Boosting, and 13) Logistic/Linear Regression. We used the default configurations provided by Sckit-learn (Pedregosa et al., 2011) for these stackers. The input to the models is the concatenation of all the base models’ predictions. We concatenated the probabilistic predictions from all the classes in the classification tasks. This sometimes produced a large dimensional input space. Finally, we include models from Dynamic Ensemble Search (DES) literature that are in the DESlib library (Cruz et al., 2020) such as KNOP (Cavalin et al., 2013), KNORAE (Ko et al., 2008) and MetaDES (Cruz et al., 2015).
@@
We train the neural networks for $10000$ update steps, with a batch size of $2048$. If the GPU memory is not enough for fitting the network because of the number of base models, or the number of classes, we decreased the batch size to $256$. Additionally, we used the Adam optimizer and a network with four layers, $32$ neurons, and a probability of keeping base models $\gamma=0.25$, or alternatively a DropOut rate $\delta=0.75$. Notice that the architecture of the ensemblers slightly varies depending on the mode (Stacking or Model Average). For the Stacking mode, we use an MLP with four layers and $32$ neurons with ReLU activations. For MA mode, we use two MLPs as in Equation 6: 1) $\textrm{MLP}_{1}$ has $3$ layers with $32$ neurons, while 2) $\textrm{MLP}_{2}$ has one layer with the same number of neurons. Although changing some of these hyperparameters might improve the performance, we keep this setup constant for all the experiments, after checking that the Neural Ensemblers perform well in a subset of the Quick-Tune metadataset (extended version).
4.3 Research Questions and Associated Experiments
Table 2: Average Normalized Error.
FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Class (AUC)
@@ 4.3 Research Questions and Associated Experiments
NE-Stack (Ours) 0.9491±0.0451 0.6331±0.0378 0.5836±0.1592 0.6104±0.3656 0.7545±0.2960 1.0440±0.3309 1.0035±0.5295
NE-MA (Ours) 0.9527±0.0402 0.6307±0.0363 0.5621±0.1483 0.8297±0.4974 0.8236±0.2240 0.9592±0.2144 0.9028±0.2157
@@
Experimental Protocol. To answer this question, we compare the neural ensembles in stacking and averaging mode to the baselines across all the metadatasets. We run every ensembling method three times for every dataset. In all the methods we use the validation data for fitting the ensemble, while we report the results on the test split. Specifically, we report the average across datasets of two metrics: negative log-likelihood (NLL) and error. For the tabular classification, we compute the ROC-AUC. As these metrics vary for every dataset, we normalize metrics by dividing them by the single-best metric. Therefore, a method with a normalized metric below one is improving on top of using the single best base model. We report the standard deviation across the experiments per dataset and highlight in bold the best method.
@@
Results. The results reported in Table 2 and Table 3 show that our proposed neural networks are competitive post-hoc ensemblers. In general, we observe that the Neural Ensemblers variants obtain either the best (in bold) or second best (italic) performance across almost all meta-datasets and metrics. Noteworthily, the greedy approach is very competitive, especially for the FTC and TR-Class. This is coherent with previous work supporting greedy ensembling as a robust method for tabular data (Erickson et al., 2020). We hypothesize that dynamic ensembling contributes partially to the strong results for the Neural Ensemblers. However, the expressivity gained is not enough, because it can lead to overfitting. To understand this, we compare to Dynamic Ensemble Selection (DES) methods. Specifically, we use KNOP, MetaDES, and KNORAE, and evaluate all methods in Scikit-learn Pipelines metadataset, as we can easily access the fitted models. We report the results of the test split in Figure 4, where we distinguish among four types of models to facilitate the reading: Neural, DES, Stacking and Constant. We can see that Neural Ensemblers are the most competitive approaches, especially on stacking mode. Additionally, we report the metrics on the validation split in Figure 8 (Appendix E), where we observe that some dynamic ensemble approaches such as Gradient Boosting (GBT), Random Forest (RF) and KNORAE exhibit overfitting, while Neural Enemblers are more robust in this sense.
Table 3: Average Normalized NLL.
FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Greedy 0.6943±0.0732 0.5785±0.1972 1.0000±0.0000 0.9025±0.2378 0.9093±0.1017 0.9665±0.0926 1.0149±0.1140
Gradient Boosting 0.7159±0.1529 1.7288±1.2623 1.2764±0.4787 1.9373±1.2839 2.6193±2.3159 1.4288±1.2083 1.0498±0.2128
DivBO 0.7695±0.1195 0.7307±0.3061 0.7125±0.3982 1.2251±1.0293 0.9430±0.2036 1.0023±0.3411 1.0247±0.1473
NE-Stack (Ours) 0.7562±0.1836 0.5278±0.2127 0.6336±0.3456 0.7486±0.6831 0.6769±0.2612 1.3268±0.7498 1.2379±0.4083
NE-MA (Ours) 0.6952±0.0730 0.5822±0.2147 0.6522±0.3131 1.0177±0.5151 0.9166±0.0936 1.0515±1.0003 0.9579±0.0777
RQ 2: Do Neural Ensemblers need a strong group of base models, i.e. found using Bayesian Optimization?
Results. We report in Table 4 the results with the two selection methods (random and DivBO) using a subset of common baselines, where we normalize using the metric of the single-best from the whole set of models. We limit the number of baselines for brevity, and the extended version includes more baselines in Table 15. We also include results in separate tables for the two selection methods (Appendix E). We can compare directly with the results in Table 3. We observe that reducing the number of base models with DivBO negatively affects the performance of the Neural Ensemblers. Surprisingly, randomly selecting the subset of base models improves the results in two metadatasets (TR-Class and NB-1000). We hypothesize that decreasing the number of base models is beneficial for these metadatasets. With over $1000$ base models available, the likelihood of identifying a preferred model and overfitting the validation data increases in these metadatasets. Naturally, decreasing the number of base models can also be detrimental for the Neural Ensemblers, as this happens for some metadatasets such as TR-Reg and QT-Micro. In contrast to the Neural Ensemblers, selecting a subset of strong models with DivBO improves the performance for some baselines such as Model Averaging (MA) or TopK ($K=25$). In other words, it works as a preprocessing method for these ensembling approaches. Overall, the results in Tables 4 and 15 demonstrate that Neural Ensemblers do not need a strong group of base models to achieve competitive results.
Table 4: Average Normalized NLL with a Subset of Base Models.
Selector FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single - 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Single DivBO 1.0000±0.0000 1.0000±0.0000 0.8707±0.3094 1.7584±2.0556 1.1846±0.2507 1.1033±0.9951 1.0039±0.0424
Random DivBO 0.9305±0.3286 0.6538±0.2123 0.9724±0.0478 1.1962±1.0189 0.9717±0.1919 1.0107±0.3431 1.0302±0.1250
Top25 DivBO 0.7617±0.1136 0.5564±0.1961 0.9762±0.0413 1.1631±0.9823 0.9431±0.2035 1.0023±0.3411 1.0247±0.1473
Quick DivBO 0.7235±0.0782 0.6137±0.1945 0.9646±0.0614 1.2427±1.1130 0.9544±0.2050 1.0014±0.3423 1.0400±0.1949
Greedy DivBO 0.7024±0.0720 0.6839±0.3003 0.9762±0.0413 1.1659±0.9789 0.9435±0.2029 1.0024±0.3410 1.0271±0.1531
MA DivBO 0.7245±0.0788 0.5712±0.2185 0.9678±0.0558 1.0559±0.7452 0.9501±0.1617 1.0068±0.4141 1.0237±0.1502
Single Random 1.0067±0.0164 1.0000±0.0000 0.9240±0.3504 1.2915±0.9952 1.1261±0.3134 1.0225±0.3353 1.1378±0.4641
Top25 Random 0.8397±0.1000 0.5848±0.1980 0.6526±0.3019 3.6553±2.7053 3.0436±2.1378 1.2599±1.5015 1.0611±0.2799
Quick Random 0.7305±0.0764 0.5958±0.1917 0.6656±0.2968 1.7769±2.1443 1.1646±0.3728 1.0797±1.0007 1.0151±0.1546
Greedy Random 0.7024±0.0720 0.5783±0.1857 0.6617±0.2839 1.6723±2.1446 0.9961±0.1290 1.0725±0.9978 1.0023±0.0961
MA Random 0.9069±0.1812 0.8677±0.2292 0.6698±0.2898 4.8593±3.1360 3.4575±2.6490 1.4759±1.9396 1.4286±1.7242
NE-Stack (Ours) DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.5040±1.9442 0.8329±0.2659 0.9729±0.3952 6.9453±3.4749
NE-MA (Ours) DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 1.1237±0.9964 0.9200±0.1966 1.0016±0.3407 1.0070±0.0977
NE-Stack (Ours) Random 0.7709±0.2204 0.7551±0.2493 0.6187±0.2950 0.8292±0.5466 0.8160±0.3852 0.9540±0.5077 4.2183±3.4808
NE-MA (Ours) Random 0.6972±0.0712 0.7911±0.2147 0.6650±0.2750 1.6877±2.1535 1.0903±0.2578 1.0674±0.9998 1.0277±0.1994
RQ 3: What is the impact of the DropOut regularization scheme?
@@
Experimental Protocol. Firstly, we do a small experiment to observe how the DropOut rate affects the weights when using Neural Ensemblers in MA mode. For this, we use QT-Micro datasets, and save each model’s weights $\theta_{m}$ and error after training the neural network. Subsequently, to understand how much the base learners DropOut helps the Neural Ensemblers, we run an ablation by trying the following values for the DropOut rate $\delta\in\{0.0,0.1,\dots,0.9\}$. We compute the average NLL for three seeds per dataset and divide this value by the one obtained for $\delta=0.0$ in the same dataset. Therefore, we realize that a specific DropOut rate is improving over the default network without regularization if the normalized NLL is below 1.
@@
Results. We show how the mean of the weights per model is related to the mean error of the models for different DropOut rates in Figure 5. When there is no DropOut some weights are close to one, i.e. they are preferred models. As we increase the value, many models with high weights decrease. If the rate is very high (e.g. $0.9$), we will have many models contributing to the ensemble, with weights different from zero. Our ablation study demonstrates that non-existing or high DropOut are detrimental to the Neural Ensembler performance in general. As shown in Figure 6, this behavior is consistent in all datasets and both modes, but TR-Reg metadataset on Stacking mode. In general, we observe that Neural Ensemblers obtain better performance when using base models’ DropOut.

### v2

@@ 4  Experiments and Results
In this section, we provide empirical evidence of the effectiveness of our approach.
@@ 4.1  Experimental Setup
In our experiments, we utilize four meta-datasets with pre-computed predictions, which allows us to simulate ensembles without the need to fit models. These meta-datasets cover diverse data modalities, including Computer Vision, Tabular Data, and Natural Language Processing. Additionally, we evaluate the method on datasets without pre-computed predictions to assess the performance of ensembling methods that rely on fitted models during the evaluation. Table 11 reports the main information related to these datasets. Particularly for Nasbench, we created 2 versions by subsampling 100 and 1000 models. The meta-dataset Finetuning Text Classifiers (FTC) (Arango et al., 2024b, ) contains the predictions of several language models to evaluate ensembling techniques on text classification tasks by finetuning language models such as GPT2 (Radford et al.,, 2019), Bert (Devlin et al.,, 2018) and Bart (Lewis et al.,, 2019). We also generate a set of fitted Scikit-Learn Pipelines on classification datasets. In this case, we stored the pipeline in memory, allowing us to evaluate our method in practical scenarios where the user has fitted models instead of predictions. We detailed information about the creation of this meta-dataset in Appendix O. Altogether, the meta-datasets comprise over 240 datasets in total. The information about each data set lies in the referred work under the column Task Information in Table 11.
Table 1: Average Normalized NLL across Metadatasets.
FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Reg
Greedy 0.6943±0.0732 0.5785±0.1972 0.7617±0.3435 0.9025±0.2378 0.9093±0.1017 0.9665±0.0926 1.0149±0.1140
Gradient Boosting 0.7159±0.1529 1.7288±1.2623 1.2575±0.4460 1.9373±1.2839 2.6193±2.3159 1.4288±1.2083 1.0498±0.2128
XGBoost 0.8292±0.1434 0.7389±0.2326 0.9092±0.5304 3.7822±3.1194 2.6119±2.3911 1.7697±1.4672 1.2580±0.4875
CatBoost 0.6887±0.0953 0.8092±0.2513 0.9512±0.5083 2.6262±2.6482 2.4145±1.8989 1.2570±1.2859 1.0454±0.1550
LightGBM 0.7973±0.1946 3.6004±2.5822 5.3943±4.7980 3.0378±2.7945 3.6860±3.2856 1.8298±1.1596 1.6250±2.1651
Akaike 0.8526±0.1403 0.5838±0.2031 0.6485±0.3166 3.1574±2.5898 2.6888±2.0620 1.0930±1.0203 1.0221±0.1793
DivBO 0.7695±0.1195 0.7307±0.3061 0.6628±0.3435 1.2251±1.0293 0.9430±0.2036 1.0023±0.3411 1.0247±0.1473
NE-Stack 0.7562±0.1836 0.5278±0.2127 0.6336±0.3456 0.7486±0.6831 0.6769±0.2612 1.3268±0.7498 1.2379±0.4083
NE-MA 0.6952±0.0730 0.5822±0.2147 0.6522±0.3131 1.0177±0.5151 0.9166±0.0936 1.0515±1.0003 0.9579±0.0777
@@
We compare the Neural Ensemblers (NE) with other common and competitive ensemble approaches. 1) Single best selects the best model according to the validation metric; 2) Random chooses randomly $N=50$ models to ensemble, 3) Top-N ensembles the best $N$ models according to the validation metric; 4) Greedy creates an ensemble with $N=50$ models by iterative selecting the one that improves the metric as proposed by previous work (Caruana et al.,, 2004); 5) Quick builds the ensemble with 50 models by adding model subsequently only if they strictly improve the metric; 6) CMAES (Purucker and Beel, 2023b, ) uses an evolutive strategy with a post-processing method for ensembling, 7) Model Average (MA) computed the sum of the predictions with constant weights as in Equation 3. We also compare to methods that perform ensemble search iteratively via Bayesian Optimization such as 8) DivBO (Shen et al.,, 2022), and 9) Ensemble Optimization (EO) (Levesque et al.,, 2016). We also report results by using common ML models as stackers, such as SVM and Random Forest. Finally, we include models from Dynamic Ensemble Search (DES) literature such as KNOP (Cavalin et al.,, 2013), KNORAE (Ko et al.,, 2008) and MetaDES (Cruz et al.,, 2015). We provide further details on the baselines setup in Appendix E.
Figure 3: Results on Scikit-Learn Pipelines.
@@
We train the neural networks for $10000$ update steps, with a batch size of $2048$. If the GPU memory is not enough for fitting the network because of the number of base models, or the number of classes, we decreased the batch size to $256$. Additionally, we used the Adam optimizer and a network with four layers, $32$ neurons, and a probability of keeping base models $\gamma=0.25$, or alternatively a DropOut rate $\delta=0.75$. Notice that the architecture of the ensemblers slightly varies depending on the mode, Stacking or Model Average (MA). For the Stacking mode, we use an MLP with four layers and $32$ neurons with ReLU activations. For MA mode, we use two MLPs as in Equation 6: 1) $\textrm{MLP}_{1}$ has $3$ layers with $32$ neurons, while 2) $\textrm{MLP}_{2}$ has one layer with the same number of neurons. Although changing some of these hyperparameters might improve the performance, we keep this setup constant for all the experiments, after checking that the Neural Ensemblers perform well in a subset of the Quick-Tune meta-dataset (extended version).
Table 2: Average Normalized Error across Metadatasets.
FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Class (AUC)
@@
NE-Stack 0.9491±0.0451 0.6331±0.0378 0.5836±0.1592 0.6104±0.3656 0.7545±0.2960 1.0440±0.3309 1.0035±0.5295
NE-MA 0.9527±0.0402 0.6307±0.0363 0.5621±0.1483 0.8297±0.4974 0.8236±0.2240 0.9592±0.2144 0.9028±0.2157
4.2  Research Questions and Associated Experiments
@@
Experimental Protocol. To answer this question, we compare the neural ensembles in stacking and averaging mode to the baselines across all the meta-datasets. We run every ensembling method three times for every dataset. In all the methods we use the validation data for fitting the ensemble, while we report the results on the test split. Specifically, we report the average across datasets of two metrics: negative log-likelihood (NLL) and classification error. For the tabular classification, we compute the ROC-AUC. As these metrics vary for every dataset, we normalize metrics by dividing them by the single-best metric. Therefore, a method with a normalized metric below one is improving on top of using the single best base model. We report the standard deviation across the experiments per dataset and highlight in bold the best method.
@@
Results. The results reported in Table 2 and Table 1 show that our proposed regularized neural networks are competitive post-hoc ensemblers. In general, we observe that the Neural Ensemblers variants obtain either the best (in bold) or second best (italic) performance across almost all meta-datasets and metrics. Noteworthily, the greedy approach is very competitive, especially for the FTC and TR-Class meta-datasets. This is coherent with previous work supporting greedy ensembling as a robust method for tabular data (Erickson et al.,, 2020). We hypothesize that dynamic ensembling contributes partially to the strong results for the Neural Ensemblers. However, the expressivity gained is not enough, because it can lead to overfitting. To understand this, we compare to Dynamic Ensemble Selection (DES) methods. Specifically, we use KNOP, MetaDES, and KNORAE, and evaluate all methods in Scikit-learn Pipelines meta-dataset, as we can easily access the fitted models. We report the results of the test split in Figure 3, where we distinguish among four types of models to facilitate the reading: Neural, DES, Stacking and Constant. We can see that Neural Ensemblers are the most competitive approaches, especially on stacking mode. Additionally, we report the metrics on the validation split in Figure 6 (Appendix F), where we observe that some dynamic ensemble approaches such as Gradient Boosting (GBT), Random Forest (RF) and KNORAE exhibit overfitting, while Neural Enemblers are more robust.
RQ 2: What is the impact of the DropOut regularization scheme?
@@
Experimental Protocol. To understand how much the base learners DropOut helps the Neural Ensemblers, we run an ablation by trying the following values for the DropOut rate $\delta\in\{0.0,0.1,\dots,0.9\}$. We compute the average NLL for three seeds per dataset and divide this value by the one obtained for $\delta=0.0$ in the same dataset. In this setup, we realize that a specific DropOut rate is improving over the default network without regularization if the normalized NLL is below 1.
@@
Results. Our ablation study demonstrates that non-existing or high DropOut are detrimental to the Neural Ensembler performance in general. As shown in Figure 4, this behavior is consistent in all datasets and both modes, but TR-Reg metadataset on Stacking mode. In general, we observe that Neural Ensemblers obtain better performance when using base models’ DropOut. Furthermore, we show how the mean weight per model is related to the mean error of the models for different DropOut rates in Figure 15 in the Appendix N.
@@
RQ 3: How sensible are the Neural Ensemblers to its hyperparameters?
Experimental Protocol. While our proposed Neural Ensembler (NE) uses MLPs with 4 layers and 32 neurons per layer by default, it is important to understand how sensitive the NE’s performance is to changes in its hyperparameters. To this end, we conduct an extensive ablation study by varying the number of layers $L\in\{1,2,3,4,5\}$ and the number of neurons per layer (hidden dimension) $H\in\{8,16,32,64,128\}$ in the stacking mode of the NE. For each configuration, we compute the Negative Log Likelihood (NLL) on every task in the metadatasets and normalize these values by the performance obtained with the default hyperparameters ($L=4$, $H=32$). This normalization allows us to compare performance changes across different tasks and metadatasets, accounting for differences in metric scales. A normalized NLL value below one indicates improved performance compared to the default setting.
Results. Our findings indicate that there is no single hyperparameter configuration that is optimal across all datasets. However, our default configuration ($L=4$, $H=32$) strikes a balance, providing robust performance across diverse tasks without the need for extensive hyperparameter tuning, and can be effectively applied in various settings without the need for dataset-specific hyperparameter optimization.
Significance on Datasets with a lot of Classes
Given the high performance between Greedy and NE-MA, we wanted to understand when the second one would obtain strong significant results. We found that NE-MA is particularly well-performing in datasets with a large number of classes. Given Table 11, we can see that four meta-datasets have a high ($>10$) number of classes. We selected these metadatasets (NB(100), NB(1000), QT-Micro, QT-Mini), and plotted the significance compared to Greedy and Random Search. The results reported in Figure 9 demonstrate that our approach is significantly better than Greedy in these metadatasets.
Additional research questions
We further discuss interesting research questions in the Appendix due to the limited space. Some of these include: i) how much the validation data size affects the Neural Ensembler (Appendix J), ii) what happens if we train the Neural Ensemblers on a merged split containing both training and validation data (Appendix K), and iii) the disadvantages of Neural Ensemblers acting on the original input space (Appendix L).
Figure 5: Ablation study of Neural Ensembler hyperparameters: number of neurons per layers (a) and number of layers (b). Normalized NLL values below one indicate improved performance over the default setting ($L=4$, $H=32$).